# ChipGPT: How far are we from natural language hardware design

### v3

@@ Abstract
Index Terms:
agile hardware development, natural language programming, program synthesis
@@ II-A Agile Hardware Design Workflow
Figure 2: Ideal Agile method productivity
Figure 3: ChipGPT productivity improvement
@@ III Automatic Chip Generation Framework
Third, to improve the quality of the $code$(raw program) generated by the GPT model, we propose an output manager $OM$.
Finally, the code generated in the third stage typically does not meet performance, power and area (PPA) requirements. To obtain an optimized final output program for PPA, we propose a search method compatible with the previous stages that ranks GPT-generated programs and identifies the version according to the $requirements$.
@@ III-A Specification Split
$eg$ represents examples within the module specifications. This information is often provided in tables where each row lists inputs and outputs within a clock cycle. ChipGPT does not take $eg$ as the prompt input, for the reason that this information is too complicated to fit the transformer model. Rather, we handle it in the post-processing step. The Output manager leverages this information to verify module correctness, reducing the need for manual human correction.
$iface$ represents the input and output ports definition.
$func$ represents the function of the module, which is the most crucial part of the ChipGPT workflow. The module function defines its core purpose and behavior.
$compose$ delineates how to compose the module, as designated when the module comprises the top level of the design.
III-B Prompt Manager $PM$
@@
[…]
@@ III-B3 Module Composition Principle
Based on this observation, we propose the ”bottom-up composition principle” to improve the quality of the generated raw program, which is similar to least-to most principle[25]. As shown in Equation 4, this principle specifies that submodule interfaces must be declared before the top module. For example, $<$ the submodule has interface $\dots$ $>$ followed by $<$ the top module consists of $\dots$ $>$. By following this sequence, when GPT generates the top module, the submodules’ ports are clearly defined. As a result, GPT can connect them appropriately. Because this principle is recursive, it can be applied to a wide range of architectural design scenarios. For example, in Fig. 9, we want to design a complex module. On the left, the add module and minus module are defined in the same prompt, where the LLM cannot output the correct implementation of the submodules. However, when applying the composition principle, the LLM can generate correct implementations of the submodules.
@@ III-C1 Correction Manager
$Machine\_correct=simulator(synthesis(raw),eg)$ (7)
@@ III-C2 Cost Manager
Figure 10: Matrix Multiply Verilog Program list, where power is measured in milliwatts ($mW$), area is measured in micrometers squared (${\mu}m^{2}$)
@@ IV-A2 Programbility measurement
$quality=raw+correct-prompt$ (9)
Where $quality$ represents code generation algorithm efficiency, $raw$ is the number of lines in the generated program, $correct$ is the number of lines modified by humans to correct the raw program, and $prompt$ is the number of prompt queries.
@@ IV-A4 Baseline
TABLE VII: PPA comparison to show the efficiency of our method under the area optimization target, where power is measured in milliwatts ($mW$), area is measured in micrometers squared (${\mu}m^{2}$) and latency measures the number of running cycles
@@
adder-multi tree ChatGPT(Baseline) 28.50 60070.4 1
@@
Figure 12: Generated code line number
Figure 13: Code quality
@@ IV-B Experimental Results
Evidence: Fig. 14 show that a 2.01$\times$ average increase in code quality is brought by the three principles in the prompt manager to our framework. They can enhance single-module generation quality and consistency for complex, large-scale designs.

### v4

@@ III Automatic Chip Generation Framework
Third, to improve the quality of the codecode(raw program) generated by the GPT model, we propose an output manager $OM$.
Finally, the code generated in the third stage typically does not meet performance, power and area (PPA) requirements. To obtain an optimized final output program for PPA, we propose a search method compatible with the previous stages that ranks GPT-generated programs and identifies the version according to the requirementsrequirements.
@@ III-A Specification Split
egeg represents examples within the module specifications. This information is often provided in tables where each row lists inputs and outputs within a clock cycle. ChipGPT does not take $eg$ as the prompt input, for the reason that this information is too complicated to fit the transformer model. Rather, we handle it in the post-processing step. The Output manager leverages this information to verify module correctness, reducing the need for manual human correction.
ifaceiface represents the input and output ports definition.
funcfunc represents the function of the module, which is the most crucial part of the ChipGPT workflow. The module function defines its core purpose and behavior.
composecompose delineates how to compose the module, as designated when the module comprises the top level of the design.
III-B Prompt Manager PMPM
@@
To improve prompt quality, we design a template-based prompt manager where prompts are interconnected within templates. We first provide a formal representation for the prompt manager. Let $S=\{(Q_{1},A_{1}),(Q_{2},A_{2}),\dots,(Q_{N},A_{N})\}$ denote a dialogue sequence with $N$ question-answer pairs. GPT takes queries as input and generates the final answer. We define archarch in Equ. 2 as a query sequence, or prompt sequence illustrated in Fig. 6. The query sequence comprises three partitions. Elements enclosed in <><> represent a sequence of prompts, while elements in $\{\}$ represent groups combined into a single prompt. The first setup prompt << act as a proficient Verilog programmer>> establishes an execution environment enabling the following queries to operate in a Verilog context. Next, for the basic modules in the specification, GPT generates $N$ modules following the interface model principle in III-B1 and post-addition principle in III-B2. Finally, there is a top module following the composecompose rule that connects these submodules from the second step. This composition adheres to the module composition principle in III-B3.
@@ III-B3 Module Composition Principle
Based on this observation, we propose the ”bottom-up composition principle” to improve the quality of the generated raw program, which is similar to least-to most principle[25]. As shown in Equation 4, this principle specifies that submodule interfaces must be declared before the top module. For example, << the submodule has interface $\dots$ >> followed by << the top module consists of $\dots$ >>. By following this sequence, when GPT generates the top module, the submodules’ ports are clearly defined. As a result, GPT can connect them appropriately. Because this principle is recursive, it can be applied to a wide range of architectural design scenarios. For example, in Fig. 9, we want to design a complex module. On the left, the add module and minus module are defined in the same prompt, where the LLM cannot output the correct implementation of the submodules. However, when applying the composition principle, the LLM can generate correct implementations of the submodules.
@@ III-C1 Correction Manager
Machine_correct=simulator(synthesis(raw),eg)Machine\_correct=simulator(synthesis(raw),eg) (7)
@@ III-C2 Cost Manager
Figure 10: Matrix Multiply Verilog Program list, where power is measured in milliwatts (mWmW), area is measured in micrometers squared (${\mu}m^{2}$)
@@ IV-A2 Programbility measurement
quality=raw+correctpromptquality=raw+correct-prompt (9)
Where qualityquality represents code generation algorithm efficiency, rawraw is the number of lines in the generated program, correctcorrect is the number of lines modified by humans to correct the raw program, and promptprompt is the number of prompt queries.
@@ IV-A4 Baseline
TABLE VII: PPA comparison to show the efficiency of our method under the area optimization target, where power is measured in milliwatts (mWmW), area is measured in micrometers squared (${\mu}m^{2}$) and latency measures the number of running cycles
@@
adder-multi
tree
ChatGPT(Baseline) 28.50 60070.4 1
@@ IV-B Experimental Results
[…]